A chemistry graduate student is given 500.mL of a 0.60M hydrocyanic acid HCN solution. Hydrocyanic acid is a weak acid with =Ka×4.910−10. What mass of NaCN should the student dissolve in the HCN solution to turn it into a buffer with pH =9.81? You may assume that the volume of the solution doesn't change when the NaCN is dissolved in it. Be sure your answer has a unit symbol, and round it to 2 significant digits.
Answer #1

HCN NACN weak acid conjugate base 0.60M ka 9 HON = 4.9 x10-10 bra = -log (ka) = -log (4.9x10-10) = 9.31 volume g solution - smol of NacN = a = molarita Xvolume 1.896 mol x 0.5% = 0.948 mol molar mang NaN = 49.00 g mol man f Nacht = molt molar mass 0.